# Efficient Length-Generalizable Attention via Causal Retrieval for Long-Context Language Modeling

### v2

@@ Abstract
Despite the success of Transformers, handling long contexts remains challenging due to the limited length generalization and quadratic complexity of self-attention. Thus Transformers often require post-training with a larger attention window, significantly increasing computational and memory costs. In this paper, we propose a novel attention mechanism based on dynamic context, Grouped Cross Attention (GCA), which can generalize to 1000 the pre-training context length while maintaining the ability to access distant information with a constant attention window size. For a given input sequence, we split it into chunks and use each chunk to retrieve top- relevant past chunks for subsequent text generation. Specifically, unlike most previous works that use an off-the-shelf retriever, our key innovation allows the retriever to learn how to retrieve past chunks that better minimize the auto-regressive loss of subsequent tokens in an end-to-end manner. Such a mechanism accommodates retrieved chunks with a fixed-size attention window to achieve long-range information access, significantly reducing computational and memory costs during training and inference. Experiments show that GCA-based models achieve near-perfect accuracy in passkey retrieval for 16M context lengths, which is the training length.
@@ 1 Introduction
Transformers (Vaswani et al., 2017), serving as the backbone of large language models (LLM), have demonstrated exceptional performance across a wide range of natural language processing tasks (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024). While Transformers excel in representational power, they still struggle to process longer contexts as expanding the attention window beyond the pre-trained size may disrupt attention distribution (Wang et al., 2024). Some methods use a sliding window mechanism (Xiao et al., 2024) with a fixed attention field to achieve extrapolation, but they fail to capture long-distance context dependencies outside the attention window. Thus, a main challenge for the attention mechanism lies in how to achieve length generalization but still maintain random access to long-range information. To mitigate the issue, most long-context LMs (Reid et al., 2024) resort to expanding the attention window (Liu et al., 2023) during post-training, significantly increasing training and inference costs due to the quadratic complexity of self-attention.
In this work, inspired by retrieval-based language models (RLM) (Asai et al., 2023), we introduce an efficient attention mechanism that achieves both length generalization and long-range information accessibility. Typically, RLMs (Rubin & Berant, 2024) divide an input sequence into chunks, retrieve relevant ones from the history for the current input, and then integrate the retrieved chunks into the decoder to predict subsequent tokens. By choosing top- chunks as a “dynamic context”, RLMs can maintain the ability to access distant information without expanding the attention window. However, most RLMs (Lewis et al., 2020; Borgeaud et al., 2022) rely on separately pre-trained retrievers, whose retrieved chunks are not necessarily helpful to the causal LMs. Although a straightforward approach is training the retriever end-to-end with the auto-regressive loss, it is rarely explored. The main difficulty is the chunk relevance scores computed by the retriever do not participate in the next token prediction, thus unable to receive gradient backpropagation from the auto-regressive loss. To overcome the difficulty, we propose Grouped Cross-Attention (GCA), a novel attention mechanism that integrates chunk-to-chunk retrieval as an inherent ability, thus the retriever can learn to retrieve past chunks that most effectively reduce the auto-regressive loss of subsequent tokens, which we refer to as causal retrieval.
@@
To fairly compare GCA with other attention mechanisms, we pre-train all models from scratch and evaluate them on tasks such as long-context language modeling, summarization, and the needle-in-a-haystack (NIAH) tests. The results demonstrate that DRT significantly outperforms all baselines with comparable pre-training costs and much lower inference costs. Notably, in the NIAH test, DRT maintains perfect accuracy on inputs up to 16M tokens. More interestingly, case studies on the arXiv-math dataset validate the causal retrieval ability in DRT, which retrieves lemmas, variants, or functions defined distantly but used in the next chunk. These findings suggest that GCA has the potential to be a fundamental component in long-context LMs.
Overall, our main contributions are:
@@
We propose a novel attention mechanism called Grouped Cross-Attention (GCA) and its hardware-aware implementation, which simultaneously achieves length generalization and long-range random-access.
@@
We conduct comprehensive experiments and demonstrate promising results of GCA. To the best of our knowledge, GCA is the first attention mechanism that can achieve perfect passkey retrieval with 16M context length, 1000 the pre-training length.
@@ 2 Related works
There are two long-context LMs closely related to ours. One of them is Retrieval-Pretrained Transformer (RPT) (Rubin & Berant, 2024). The key difference between DRT and RPT is the training approach of the retriever. During data-preparation, for each chunk, RPT picks relevant past chunks by using BM25 (Robertson & Zaragoza, 2009), concatenates them with the current chunk, and evaluates them by a reference LM like Pythia 1.4B (Biderman et al., 2023). The past chunks that increase the probability of the next chunk are identified as ‘gold chunks’ to train RPT’s retriever. This process requires that the reference LM be pre-trained on the target dataset beforehand, significantly limiting scalability and flexibility. Landmark Attention (LA) (Mohtashami & Jaggi, 2023) is another close work. LA is pre-trained with short contexts but capable of handling long contexts during inference. It addresses long-context language modeling by modifying self-attention KV Cache. During inference, each token, at each layer, selects top- chunks based on token-to-chunk attention scores and appends their key and value vectors to the current KV cache of self-attention. The token-to-chunk attention scores are trained in an end-to-end manner with a grouped softmax technique. However, it has to perform top- chunk selection per token, per layer, which incurs significant extra costs during inference. Moreover, it fails to extrapolate on longer context length. Our method combines the chunk-retrieval and grouped softmax ideas, resolving the aforementioned issues while balancing training efficiency and inference performance.
@@
Long-Context Language Modeling.
Various methods have been proposed to improve long-context language modeling. A straightforward approach is extending the attention window and conducting post-training on a longer context (Bai et al., 2024; Reid et al., 2024; Qwen et al., 2025). However, it comes with a significant additional training cost and still has length generalization issues. One line of research is introducing memorization to Transformers via recurrence. Many works (Dai et al., 2019; Burtsev & Sapunov, 2020; Martins et al., 2022; Hutchins et al., 2022) compress past information into fixed-sized vectors. However, these methods often sacrifice the flexibility to attend to arbitrary past tokens. Meanwhile, other works focus on maintaining random-access flexibility of attention. Memorizing Transformers (Wu et al., 2022) appends retrieved past keys and values to the current attention segment via -NN search, but they do not back-propagate gradients to them. CEPE (Yen et al., 2024) parallelly encode long-conext by chunks and fuses them into the decoder via cross-attention. During the training process, the decoder parameters are fixed, and only the encoder is adjusted. Recently, state-space models (Gu & Dao, 2023; Dao & Gu, 2024) and RNN models (Beck et al., 2024) provide new architecture alternatives, with comparable performance to Transformers but much lower cost for inference. A notable distinction of our work is able to achieve perfect past information retrieval for input length 1000 the pre-training length.
@@ 3 Methodology
A typical architecture of RLMs (Borgeaud et al., 2022; Yen et al., 2024; Rubin & Berant, 2024) appends Chunked Cross-Attention (CCA) after self-attention to fuse information from retrieved chunks, in which a retriever is merely used to pick relevant chunks. Our approach makes retriever learnable by replacing CCA with GCA, which represents the key innovation of this work. The novelty of GCA lies in using relevance scores to fuse information from retrieved chunks for LM decoders, enabling the retriever to adaptively learn to select the best past chunks for predicting subsequent tokens, guided by the auto-regressive loss. This section details the model architecture, training, and inference.
@@ 3.1 Model Architecture
Here, represents the landmark representation output by the previous decoder layer of the -th group, which accumulates all information from groups 1 to . This enables to perform multiple retrievals based on retrieved information fused by previous layers. represents the causal relevance score of to . contains the indices of past chunks with top- relevance scores. The retrieved chunks are shared among the subsequent layers within the same group. The upper layers apply GCA to fuse retrieved information into the decoder.
@@
Since is shared across layers, we use the same K, V liner transformations across layers to compact model parameters and reduce memory footprint. For each head , we have defined as:
Here, is per head dimension, and are linear transformations per head, where varies across layers and are layer-shared. We use softmax-off-by-one to allow tokens in the current chunk to ignore any retrieved tokens. The final CA outputs are concatenated vectors from all heads. It is worth noting that GCA is easy to integrate with FlashAttention-2, as detailed in the pseudo-code in Appendix C.
@@ 4.1.1 Datasets
PG19 (Rae et al., 2020) is a language modeling benchmark widely used to evaluate the long-range text understanding capabilities of models. It includes a collection of full-length books from Project Gutenberg across a wide range of genres, styles, and topics. The dataset is particularly useful for evaluating models’ abilities to capture dependencies and context over long sequences of text.
@@ 4.1.2 Models
Since the official implementation of Block-Recurrent Transformer is also based on JAX, we utilized a PyTorch implementation 111https://github.com/lucidrains/block-recurrent-transformer-pytorch to ensure all baselines are running with the same framework.
@@
| Models | Retrieval | Single NIAH (RULER) | |||||||
| 4K | 8K | 16K | 32K | 64K | 128K | 256K | 16M | ||
@@
| Models | XSum | CNN/DailyMail | Variable Tracking (RULER) | |||||||||
| R-1 | R-2 | R-L | R-1 | R-2 | R-L | 1K | 4K | 16K | 64K | 256K | 4M | |
@@
4.2 Long-Context Language Modeling
In this section, we evaluate DRT against baselines in long-context language modeling on PG19 and arXiv-math, and report their respective perplexities. All models are pre-trained with the same attention field and a 16K context by default, except for baselines that cannot efficiently pre-train on long contexts. To ensure fairness, we adjusted these baselines. Detailed hyper-parameters are provided in Appendix A.
@@ 4.3 Downstream task evaluation & Efficiency Analysis
In this section, we fine-tune all baselines under same configurations and evaluate them against downstream tasks including summarization (Nallapati et al., 2016; Narayan et al., 2018), NIAH tests proposed in RULER. Then we analyze the inference cost, the relationship between training time and context length, and the extrapolation capability of DRT. In the inference cost analysis, We skip RPT because it has a similar cost to DRT. In the extrapolation experiments, we utilize the pre-trained models described in § 4.2 to assess their perplexity on longer inputs.
@@ Results.
From Figure 3(a), it can be observed that the total training time increases linearly with the increment of the sequence length. In the extrapolation experiments, both BaseLM and DRT perform well as shown in Figure 3(b). However, Landmark Attn fails to extrapolate at longer eval lengths. While both methods use a fixed attention window and dynamically select past chunks to fill it, why can GCA extrapolate over 1000 times with stable PPL and perfect passkey retrieval, but Landmark Attn fails? We speculate a major reason lies in the different objectives for optimizing retrievers. Landmark Attn optimizes for predicting the next token based on retrieved chunks, aiming for maximizing where denotes retrieved chunks. In contrast, GCA retrieves past chunks for the next chunk prediction, aiming to maximize . A single token offers limited information, which may result in false positive retrievals thus leading to inefficacy in larger retrieval spaces. While chunks provide richer semantic information, enabling generalization to larger search spaces.
@@ 4.4 Case Studies
By analyzing DRT’s retrieval results on the arXiv-math dataset, we find some intriguing cases. A case is given in Figure 4. When retrieving past chunks, the results not only include the definition of prepositions referenced in the current chunk but also lemmas to be used in the next chunk. This validates the idea of causal retrieval, allowing us to not only retrieve semantically similar content but also information that better predicts the next chunk. More case studies can be found in Appendix D.
5 Conclusion
@@ Appendix A Hyper-parameters
We employ a Llama-like architecture (Touvron et al., 2023) featuring a 12-layer, decoder-only transformer with 12 heads per layer (64 dimensions each), an embedding dimension of 768, and an FFN size of 2048 for the 128M model, and a 24-layer transformer with 16 heads per layer (64 dimensions each), and an FFN size of 2816 for the 350M model. Training utilizes the AdamW optimizer (Loshchilov & Hutter, 2019) with and , and a weight decay factor of 0.001. We used base learning rate for all our experiments with a warmup stage that was 2% of the whole training and applied a cosine scheduler with the final learning rate being . We used GPT-2’s (Radford et al., 2019) tokenizer. We used mixed-precision training with bfloat16 over at 8 Nvidia A100 GPUs. We train all models with an effective batch size of tokens for 60K steps resulting in a total training budget of 32.2 billion tokens. We train Base LM, RPT and DRT on each dataset with a context length of 16K tokens. Due to Landmark Attention not supporting sliding-window attention at the training stage, the model is pre-trained with full self-attention with a context length of 768. Due to Block Recurrent Transformer cannot be fully paralleled, which takes wall-clock training time with 16K context length, we pre-train it with a context length of 4K.
Appendix B Results for larger models
Appendix C Hardware-aware GCA psuedo-code
Appendix D More case studies

### v4

@@ Abstract
Despite the success of Transformers, handling longer contexts remains challenging due to the limited length generalization and quadratic complexity of self-attention. Transformers often require post-training with a larger attention window, significantly increasing computational and memory costs. In this paper, we propose a novel attention mechanism based on dynamic context, Grouped Cross Attention (GCA), which can generalize to 1000 the pre-training context length while maintaining the ability to access distant information with a constant attention window size. For a given input sequence, we split it into chunks and use each chunk to retrieve top- relevant past chunks for subsequent text generation. Specifically, unlike most previous works that use an off-the-shelf retriever, our key innovation allows the retriever to learn how to retrieve past chunks that better minimize the auto-regressive loss of subsequent tokens in an end-to-end manner. Such a mechanism accommodates retrieved chunks with a fixed-size attention window to achieve long-range information access, significantly reducing computational and memory costs during training and inference. Experiments show that GCA-based models achieve near-perfect accuracy in passkey retrieval for 16M context lengths, which is the training length.
@@ 1 Introduction
Transformers (Vaswani et al., 2017), serving as the backbone of large language models (LLM), have demonstrated exceptional performance across a wide range of natural language processing tasks (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024). While Transformers excel in representational power, they still struggle to process longer contexts as expanding the attention window beyond the pre-trained size may disrupt attention distribution (Wang et al., 2024). Some methods use a sliding window mechanism (Xiao et al., 2024) with a fixed attention field to achieve extrapolation, but they fail to capture long-distance context dependencies outside the attention window. Thus, a main challenge for the attention mechanism lies in how to achieve length generalization but still maintain random access to long-range information. To mitigate the issue, most long-range LMs (Reid et al., 2024) resort to expanding the attention window (Liu et al., 2023) during post-training, significantly increasing training and inference costs due to the quadratic complexity of self-attention.
In this work, inspired by retrieval-based language models (RLM) (Asai et al., 2023), we introduce an efficient attention mechanism that achieves both length generalization and long-range information accessibility. Typically, RLMs (Rubin & Berant, 2024) divide an input sequence into chunks, retrieve relevant ones from the history for the current input, and then integrate the retrieved chunks into the decoder to predict subsequent tokens. By choosing top- chunks as a “dynamic context”, RLMs can maintain the ability to access distant information without expanding the attention window. However, most RLMs (Lewis et al., 2020; Borgeaud et al., 2022) rely on separately pre-trained retrievers, whose retrieved chunks are not necessarily helpful to the causal LMs. Although a straightforward approach is training the retriever end-to-end with the auto-regressive loss, it is rarely explored. The main difficulty is that the relevance scores computed by the retriever do not participate in the next token prediction, thus cannot receive gradient backpropagation from the auto-regressive loss. To overcome the difficulty, we propose Grouped Cross-Attention (GCA), a novel attention mechanism which integrates chunk-to-chunk retrieval as an inherent ability, thus the retriever can learn to retrieve past chunks that most effectively reduce the auto-regressive loss of subsequent tokens, which we refer to as causal retrieval.
@@
To fairly compare GCA with other attention mechanisms, we pre-train all models from scratch and evaluate them on tasks such as long-range language modeling, summarization, and the needle-in-a-haystack (NIAH) tests. The results demonstrate that DRT significantly outperforms all baselines with comparable pre-training costs and much lower inference costs. Notably, in the NIAH test, DRT maintains perfect accuracy on inputs up to 16M tokens. More interestingly, case studies on the arXiv-math dataset validate the causal retrieval ability in DRT, which retrieves lemmas, variants, or functions defined distantly but used in the next chunk. These findings suggest that GCA has the potential to be a fundamental component in long-range LMs. Overall, our main contributions are:
@@
We propose a novel attention mechanism called Grouped Cross-Attention (GCA) and its hardware-aware implementation, which simultaneously achieves length generalization and long-range random-access.111The code is released at https://github.com/ant-research/long-context-modeling
@@
We conduct comprehensive experiments and demonstrate the promising results of GCA. To the best of our knowledge, GCA is the first attention mechanism that can achieve perfect passkey retrieval with 16M context length, 1000 the pre-training length.
@@ 2 Related works
There are two long-range LMs closely related to ours. One of them is Retrieval-Pretrained Transformer (RPT) (Rubin & Berant, 2024). The key difference between DRT and RPT is the training approach of the retriever. During data-preparation, for each chunk, RPT picks relevant past chunks by using BM25 (Robertson & Zaragoza, 2009), concatenates them with the current chunk, and evaluates them by a reference LM like Pythia 1.4B (Biderman et al., 2023). The past chunks that increase the probability of the next chunk are identified as ‘gold chunks’ to train RPT’s retriever. This process requires that the reference LM be pre-trained on the target dataset beforehand, significantly limiting scalability and flexibility. In contrast, GCA integrate retrieval into the attention mechanism thus can be pre-trained jointly end-to-end without reliance on any external data or models. Landmark Attention (LA) (Mohtashami & Jaggi, 2023) is another close work. LA is pre-trained with short contexts but capable of handling long contexts during inference. It addresses long-range language modeling by modifying self-attention KV Cache. During inference, each token, at each layer, selects top- chunks based on token-to-chunk attention scores and appends their key and value vectors to the current KV cache of self-attention. The token-to-chunk attention scores are trained in an end-to-end manner with a grouped softmax technique. However, it has to perform top- chunk selection per token, per layer, which incurs significant extra costs during inference. Moreover, it fails to extrapolate on longer context length. Our method combines the chunk-retrieval and grouped softmax ideas, resolving the aforementioned issues while balancing training efficiency and inference performance.
@@
Long-Range Language Modeling.
Various methods have been proposed to improve long-range language modeling. A straightforward approach is extending the attention window and conducting post-training on a longer context (Bai et al., 2024; Reid et al., 2024; Qwen et al., 2025). However, it comes with a significant additional training cost and still has length generalization issues. One line of research is introducing memorization to Transformers via recurrence. Many works (Dai et al., 2019; Burtsev & Sapunov, 2020; Martins et al., 2022; Hutchins et al., 2022; Munkhdalai et al., 2024) compress past information into fixed-sized vectors. However, these methods often sacrifice the flexibility to attend to arbitrary past tokens. Meanwhile, other works focus on maintaining random-access flexibility of attention. Memorizing Transformers (Wu et al., 2022) appends retrieved past keys and values to the current attention segment via -NN search, but they do not back-propagate gradients to them. CEPE (Yen et al., 2024) parallelly encode long-conext by chunks and fuses them into the decoder via cross-attention. During the training process, the decoder parameters are fixed, and only the encoder is adjusted. Recently, state-space models (Gu & Dao, 2023; Dao & Gu, 2024), RNN models (Beck et al., 2024) and their variants (Nunez et al., 2024) provide new architecture alternatives, with comparable performance to Transformers but much lower cost for inference. A notable distinction of our work is able to achieve perfect past information retrieval for input length 1000 the pre-training length.
@@ 3 Methodology
A typical architecture of RLMs (Borgeaud et al., 2022; Yen et al., 2024; Rubin & Berant, 2024) appends Chunked Cross-Attention (CCA) after self-attention to fuse information from retrieved chunks, in which a retriever is merely used to pick relevant chunks. Our approach makes the retriever learnable by replacing CCA with GCA, which represents the key innovation of this work. The novelty of GCA lies in using relevance scores to fuse information from retrieved chunks for LM decoders, enabling the retriever to adaptively learn to select the best past chunks for predicting subsequent tokens, guided by the auto-regressive loss. This section details the model architecture, training, and inference.
@@ 3.1 Model Architecture
Here, represents the landmark representation output by the previous decoder layer of the -th group, which accumulates all information from groups 1 to . This enables to perform multiple retrievals based on retrieved information fused by previous layers. apply linear transformations on their respective inputs. represents the causal relevance score of to . contains the indices of past chunks with top- relevance scores. The retrieved chunks are shared among the subsequent layers within the same group. The upper layers apply GCA to fuse retrieved information into the decoder.
@@
Since is shared across layers, we use the same K, V linear transformations across layers to compact model parameters and reduce memory footprint. For each head , we have defined as:
Here, is per-head dimension, and are linear transformations per head, where varies across layers and are layer-shared. We use softmax-off-by-one to allow tokens in the current chunk to ignore any retrieved tokens. The final CA outputs are concatenated vectors from all heads. It is worth noting that GCA is easy to integrate with FlashAttention-2, as detailed in the pseudo-code in Appendix B.
@@ 4.1.1 Datasets
PG19 (Rae et al., 2020) is a language modeling benchmark widely used to evaluate long-range text understanding capabilities of models. It includes a collection of full-length books from Project Gutenberg across a wide range of genres, styles, and topics. The dataset is particularly useful for evaluating models’ abilities to capture dependencies and context over long sequences of text.
@@ 4.1.2 Models
Since the official implementation of Block-Recurrent Transformer is also based on JAX, we utilized a PyTorch implementation 222https://github.com/lucidrains/block-recurrent-transformer-pytorch to ensure all baselines are running with the same framework.
@@
| Models | Retrieval | Single NIAH | |||||||
| 128M | 4K | 8K | 16K | 32K | 64K | 128K | 256K | 16M | |
@@
| Models | XSum | CNN/DailyMail | 2-hop NIAH | |||||||||
| 128M | R-1 | R-2 | R-L | R-1 | R-2 | R-L | 1K | 4K | 16K | 64K | 256K | 4M |
@@
4.2 Long-range Language Modeling
In this section, we evaluate DRT against baselines in long-range language modeling on PG19 and arXiv-math, and report their respective perplexities. All models are pre-trained with the same attention field and a 16K context by default, except for baselines that cannot efficiently pre-train on long contexts. To ensure fairness, we adjusted these baselines. Detailed hyper-parameters are provided in Appendix A.
@@ Results.
| Model | Throughput | Memory | Throughput | Memory | Throughput | Memory | Throughput | Memory |
| #params.=350M, ctx-len=32K | #params.=760M, ctx-len=16K | #params.=1.5B, ctx-len=8K | #params.=3B, ctx-len=4K | |||||
| Transformer | 50G | 50G | 50G | 72G | ||||
| Transformer | 50G | 50G | 72G | 72G | ||||
| BRT | 59G | 56G | 59G | — | OOM | |||
| RPT | 57G | 59G | 64G | 80G | ||||
| DRT | 56G | 59G | 63G | 80G | ||||
| DRT | 57G | 59G | 63G | 81G | ||||
@@ 4.3 Downstream task evaluation & Efficiency Analysis
In this section, we fine-tune all baselines under the same configurations and evaluate them against downstream tasks, including summarization (Nallapati et al., 2016; Narayan et al., 2018), NIAH tests akin to RULER. The details for the NIAH tests are described in Appendix D. Then we analyze the inference cost, the relationship between training time and context length, the training throughput, and the extrapolation capability of DRT. In the inference cost analysis, We skip RPT because it has a similar cost to DRT. In the extrapolation experiments, we utilize the pre-trained models described in § 4.2 to assess their perplexity on longer inputs.
@@ Results.
From Figure 3(a), it can be observed that the total training time increases linearly with the increment of the sequence length. In the extrapolation experiments, both BaseLM and DRT perform well as shown in Figure 3(b). However, Landmark Attn fails to extrapolate at longer eval lengths. While both methods use a fixed attention window and dynamically select past chunks to fill it, why can GCA extrapolate over 1000 times with stable PPL and perfect passkey retrieval, while Landmark Attn fails? We speculate a major reason lies in the different objectives for optimizing retrievers. Landmark Attn optimizes for predicting the next token based on retrieved chunks, aiming for maximizing where denotes retrieved chunks. In contrast, GCA retrieves past chunks for the next chunk prediction, aiming to maximize . A single token offers limited information, which may result in false positive retrievals thus leading to inefficacy in larger retrieval spaces. While chunks provide richer semantic information, enabling generalization to larger search spaces.
@@ 4.4 Case Studies
By analyzing DRT’s retrieval results on the arXiv-math dataset, we find some intriguing cases. A case is given in Figure 4. When retrieving past chunks, the results not only include the definition of prepositions referenced in the current chunk but also lemmas to be used in the next chunk. This validates the idea of causal retrieval, allowing us to not only retrieve semantically similar content but also information that better predicts the next chunk. More case studies can be found in Appendix C.
5 Conclusion & Future works
@@
Impact Statement
This paper presents a novel attention mechanism that aims to address the length generalization issue of LLMs. There are many potential societal consequences of our work associated with LLMs. Beyond LLMs, we feel no other consequences must be highlighted here.
@@ Appendix A Hyper-parameters
We employ a Llama-like architecture (Touvron et al., 2023) featuring a 12-layer, decoder-only transformer with 12 heads per layer (64 dimensions each), an embedding dimension of 768, and an FFN size of 2048. Training utilizes the AdamW optimizer (Loshchilov & Hutter, 2019) with and , and a weight decay factor of 0.001. We used base learning rate for all our experiments with a warmup stage that was 2% of the whole training and applied a cosine scheduler with final learning rate being . We used GPT-2’s (Radford et al., 2019) tokenizer. We used mixed-precision training with bfloat16 over at 8 Nvidia A100 GPUs. We train all models with an effective batch size of tokens for 60K steps resulting in a total training budget of 32.2 billion tokens. We train Base LM, RPT and DRT on each dataset with a context length of 16K tokens. Due to Landmark Attention doesn’t support sliding-window attention, the model is pre-trianed with full self-attention with a context length of 768. Due to Block Recurrent Transformer cannot be fully paralleled, which takes wall-clock training time with 16K context length, we pre-train it with a context length of 4K.
Appendix B Hardware-aware GCA psuedo-code
Appendix C More case studies
Appendix D The details for the NIAH test
In all evaluations conducted for the NIAH tests, we fine-tune all models using checkpoints derived from PG19, employing the same set of synthetic data. The number of fine-tuning steps is set to one-tenth of the total steps used during pre-training, while all other hyperparameters are kept constant. Examples of the synthetic data utilized for each task are presented in the table 5. Specifically, we pad the input tokens to ensure that the landmark token can be inserted before “is” in the question.
| Task | Example | ||||||||